How many grams of nitrogen do I have if a sample of the gas occupies a space of 4.1 liters at a pressure of 1.2 atm, and a temperature of 800 K.
1
Expert's answer
2020-04-14T01:43:34-0400

Nitrogen gas: N2;

Molar mass of N2 = Mr(N2) = 2*Ar(N) = 2*14 = 28 g/mol;

gas pressure: P = 1.2 atm;

gas volume: V = 4.1 L;

temperature: T = 800 K;

gas constant: R=0.082 L atm K−1 mol−1.


Solution:

Let's use the gas law, PV = nRT.

Rearrange the equation to isolate n and solve.

n = (PV) / (RT).

n = (1.2 * 4.1) / (0.082 * 800) = (0.075 moles).

Moles of N2 = n(N2) = 0.075 moles


Molar mass = grams / moles.

m(N2) = Mr(N2) * n(N2) = (28 g/mol) * (0.075 moles) = 2.1 g.

m(N2) = 2.1 g.


Answer: 2.1 grams of nitrogen (N2).